OpenStudy (anonymous):

y-hat=1.3+0.73x. what is the residual for (4,7)

jimthompson5910 (jim_thompson5910):

What is y-hat equal to when x = 4?

OpenStudy (anonymous):

4.22

jimthompson5910 (jim_thompson5910):

To compute the residual, simply subtract the observed-predicted So, y - (yhat) = 7 - 4.22 = 2.78 making the residual to be 2.78